OpenStudy (anonymous):

which of the following circumstances leads to a loss of internal energy of a system? A. 600 J of heat is removed from a system that has 700 J of work done on it B. 600 J of heat is removed from a system that has 600 J of work done on it C. 100 J of heat is added to a systsem that has 600 J of work done of it D. 100 J of heat is added to a system that does 600 J of work

OpenStudy (michele_laino):

here we have to apply this formula, namely the first principle of thermodynamics: \[Q = \Delta U + W\]

OpenStudy (anonymous):

ok! how do we do that?

OpenStudy (michele_laino):

we have to pick the case for which \Delta U < 0

OpenStudy (anonymous):

ok! not sure which that would be :(

OpenStudy (michele_laino):

for example, I consider the case A: there we have: Q= -600 and W= -700 so substituting, we get: \[\begin{gathered} - 600 = \Delta U - 700 \hfill \\ \Delta U = 100Joules \hfill \\ \end{gathered} \] so the answer can not be case A

OpenStudy (anonymous):

ohh i see so should it be choice C? :/

OpenStudy (michele_laino):

case B: Q=-600, W= -600, so \Delta U=0 so it is not case B

OpenStudy (anonymous):

yes:)

OpenStudy (michele_laino):

case C;: Q= +100, W= -600 so: \Delta U= 700 so it is not case C

OpenStudy (anonymous):

ohh oops :P so it is choice D?

OpenStudy (michele_laino):

yes! since case D: Q= 100, and W= 600, so: \Delta U = 100-600=-500 which is negative

OpenStudy (anonymous):

ooh yay!! thank you!!

OpenStudy (michele_laino):

thank you!! :)